# IDNet: A Novel Dataset for Identity Document Analysis and Fraud Detection

### v1

@@ Abstract
Effective fraud detection and analysis of government-issued identity documents, such as passports, driver’s licenses, and identity cards, are essential in thwarting identity theft and bolstering security on online platforms. The training of accurate fraud detection and analysis tools depends on the availability of extensive identity document datasets. However, current publicly available benchmark datasets for identity document analysis, including MIDV-500, MIDV-2020, and FMIDV, fall short in several respects: they offer a limited number of samples, cover insufficient varieties of fraud patterns, and seldom include alterations in critical personal identifying fields like portrait images, limiting their utility in training models capable of detecting realistic frauds while preserving privacy. In response to these shortcomings, our research introduces a new benchmark dataset, IDNet, designed to advance privacy-preserving fraud detection efforts. The IDNet dataset comprises $837,060$ images of synthetically generated identity documents, totaling approximately $490$ gigabytes, categorized into $20$ types from $10$ U.S. states and $10$ European countries. We evaluate the utility and present use cases of the dataset, illustrating how it can aid in training privacy-preserving fraud detection methods, facilitating the generation of camera and video capturing of identity documents, and testing schema unification and other identity document management functionalities.
@@ 3 The IDNet Benchmark Dataset
Leveraging a cutting-edge AI-assisted pipeline, outlined in Sec. 4, we have developed an identity document dataset, named IDNet, which is entirely synthetically generated and devoid of any private information. This dataset encompasses a total of $837,060$ identity documents, spanning across $20$ different document types. For each document type, there are $5,979$ unique document samples. Each sample comprises one authentic copy alongside four fraudulent variations, which include face morphing, portrait substitution, text alteration, and a combination of these fraudulent techniques as detailed in Sec. 4.4. IDNet stands as the largest publicly accessible identity document dataset to date, as depicted in Fig. 1. Detailed statistics and Zenodo URLs for each document type within IDNet are systematically presented in Tab. 1.
@@ 5.5 Utility of the Text-field Fraud Replacement Dataset in IDNet
To evaluate the synthetic fraud data generated from the text-field fraud replacement pattern, we focus on the Inpaint and Rewrite pattern and the Crop and Replace pattern, since they were also used in SIDTD [21] and the utility of the fraud detection models trained on IDNet can be evaluated on SIDTD. We selected two European identification documents, the Finland ID and the Greece Passport, for evaluation, as both document templates have been used in IDNet and SIDTD [21]. However, for each fraud pattern and each document type, our IDNet dataset contains $5979$ positive samples (w/ fraud) and $5979$ negative samples (w/o fraud), while SIDTD only contains $100$ positive samples and $1078$ negative samples for the ’Inpaint and Rewrite’ pattern, and $100$ positive samples and $144$ negative samples for the ’Crop and Replace’ pattern.
@@
Table 9: Performance comparison on Finland ID.
Dataset used for Training Dataset used for Testing
IDNet SIDTD
IDNet 1 0.941
SIDTD 0.672 1
Table 10: Performance comparison on Greece Passport.
Dataset used for Training Dataset used for Testing
IDNet SIDTD
IDNet 1 0.977
SIDTD 0.677 0.995
@@ 7 Conclusions
In this study, we introduce IDNet, a vast and comprehensive benchmark dataset comprising $597,900$ synthetic identity cards across $20$ distinct types, aimed at facilitating research in identity document analysis and privacy-enhanced fraud detection. Utilizing recent breakthroughs in artificial intelligence and large language models (LLMs), we developed a cost-effective methodology for generating representative document templates and identities. Additionally, we meticulously crafted a set of representative fraud patterns—encompassing face morphing, portrait substitution, text field alterations, and combinations thereof—that are intricately designed to intersect with personal identifier information. This intersection presents significant challenges for the development of privacy-preserving fraud detection mechanisms, as evidenced by our detailed evaluations.

### v2

@@ Abstract
Effective fraud detection and analysis of government-issued identity documents, such as passports, driver’s licenses, and identity cards, are essential in thwarting identity theft and bolstering security on online platforms. The training of accurate fraud detection and analysis tools depends on the availability of extensive identity document datasets. However, current publicly available benchmark datasets for identity document analysis, including MIDV-500, MIDV-2020, and FMIDV, fall short in several aspects: they offer a limited number of samples, cover insufficient varieties of fraud patterns, and seldom include alterations in critical personal identifying fields like portrait images, limiting their utility in training models capable of detecting realistic frauds while preserving privacy. In response to these shortcomings, our research introduces a new benchmark dataset, IDNet, designed to advance privacy-preserving fraud detection efforts. The IDNet dataset comprises $837,060$ images of synthetically generated identity documents, totaling approximately $490$ gigabytes, categorized into $20$ types from $10$ U.S. states and $10$ European countries. We evaluate the utility and present use cases of the dataset, illustrating how it can aid in training privacy-preserving fraud detection methods, facilitating the generation of camera and video capturing of identity documents, and testing schema unification and other identity document management functionalities.
@@ 3 The IDNet Benchmark Dataset
Leveraging a cutting-edge AI-assisted pipeline, outlined in Sec. 4, we have developed an identity document dataset, named IDNet, which is entirely synthetically generated and devoid of any private information. This dataset encompasses a total of $837,060$ identity documents, spanning across $20$ different document types. For each document type, there are $5,979$ unique document samples. Each sample comprises one authentic copy alongside six fraudulent variations, which include face morphing, portrait substitution, text alteration, a combination of these fraudulent techniques, and two fraud patterns (inpaint-and-rewrite, and crop-and-move) leveraged from the SIDTD benchmark, as detailed in Sec. 4.4. IDNet stands as the largest publicly accessible identity document dataset to date, as depicted in Fig. 1. Detailed statistics and Zenodo URLs for each document type within IDNet are systematically presented in Tab. 1.
@@ 5.5 Utility of the Text-field Fraud Replacement Dataset in IDNet
Table 9: Utility (detection accuracy) of the models trained on IDNet and SIDTD for detecting inpaint-and-rewrite and copy-and-move frauds for Finland ID cards.
Training Dataset Models Test Datasets
IDNet-Finland SIDTD-Finland
IDNet-Finland EfficientNet 100.0 100.0
ResNet18 100.0 100.0
ResNet50 100.0 100.0
SIDTD-Finland EfficientNet 67.2 100.0
ResNet18 67.2 100.0
ResNet50 67.2 100.0
Table 10: Utility (detection accuracy) of the models trained on IDNet and SIDTD for detecting inpaint-and-rewrite and copy-and-move frauds for Greece passports.
Training Dataset Models Test Datasets
IDNet-Greece SIDTD-Greece
IDNet-Greece EfficientNet 100.0 100.0
ResNet18 100.0 96.8
ResNet50 100.0 99.1
SIDTD-Greece EfficientNet 67.7 100.0
ResNet18 67.7 100.0
ResNet50 67.7 100.0
Detection of Inpaint-and-Rewrite, and Crop-and-Move. To evaluate the synthetic fraud data generated from the text-field fraud replacement pattern, we focus on the inpaint-and-rewrite pattern and the crop-and-replace pattern, since they were also used in SIDTD [21] and the utility of the fraud detection models trained on IDNet can be evaluated on SIDTD. We selected two European identification documents, the Finland ID and the Greece passport, for evaluation, as both document templates have been used in IDNet and SIDTD [21]. (The results on identity documents from other European countries shared by the two datasets are similar.) For each fraud pattern and each document type, our IDNet dataset contains $5979$ positive samples (w/ fraud) and $5979$ negative samples (w/o fraud). However, SIDTD only contains $100$ negative samples for each of $10$ types, and $1078$ and $144$ positive samples of the inpaint-and-rewrite and the crop-and-replace patterns for all $10$ types in total, respectively. Particularly, there are $112$ negative samples of the inpaint and rewrite fraud and $10$ samples of the copy-and-move fraud for the Finland ID card, and $109$ negative samples of the inpaint and rewrite fraud and $12$ samples of the copy-and-move fraud, for the Greece passport.
@@ 7 Conclusions
In this study, we introduce IDNet, a vast and comprehensive benchmark dataset comprising $837,060$ synthetic identity cards across $20$ distinct types, aimed at facilitating research in identity document analysis and privacy-enhanced fraud detection. Utilizing recent breakthroughs in artificial intelligence and large language models (LLMs), we developed a cost-effective methodology for generating representative document templates and identities. Additionally, we meticulously crafted a set of representative fraud patterns—encompassing face morphing, portrait substitution, text field alterations, and combinations thereof—that are intricately designed to intersect with personal identifier information. This intersection presents significant challenges for the development of privacy-preserving fraud detection mechanisms, as evidenced by our detailed evaluations.